# Anomalous dynamical scaling in interacting anyonic chains

### 16 Mar 2026 v1

@@ Abstract
Particle statistics impose fundamental constraints on nonequilibrium quantum dynamics, yet it remains an open question whether fractional statistics can lead to emergent universal dynamical scaling beyond the conventional Bose-Fermi paradigm. Here we investigate the far-from-equilibrium many-body relaxation of anyons in a one-dimensional lattice and uncover an unconventional yet universal scaling behavior governed by fractional statistics. Based on large-scale numerical simulations and scaling analysis, we identify a distinct separation between particle transport and information spreading: density correlations spread superdiffusively, whereas entanglement entropy grows ballistically. The anomalous particle dynamics can be interpreted intuitively from the statistical-phase-induced quantum interference, which suppresses coherent holon-doublon propagation. In contrast, the entanglement growth turns out to be dominated by its configurational component, which propagates ballistically. Our results establish anyonic statistics as a distinct source of universal nonequilibrium dynamics beyond bosons and fermions, with direct relevance to current quantum simulation experiments.
@@
Introduction—Fractional statistics provides a fundamental extension of quantum many-body physics beyond the conventional Bose-Fermi framework [1, 2, 3, 4, 5]. Its physical relevance has been established through the observation of exotic quantum phenomena, including the fractional quantum Hall effect [6, 7, 8, 9] and topological spin liquids [10, 11, 12], and has stimulated extensive interest in topological quantum computation [13, 14, 15, 16, 17, 18, 19]. While anyonic statistics are intrinsically associated with two-dimensional systems, anyonic behavior can also emerge in one dimension (1D). For instance, fractional exclusion statistics were first revealed through the generalized Pauli principle in the Haldane-Shastry model [20, 21]. This has triggered extensive studies of 1D anyonic models both in the continuum [22, 23, 24, 25] and on discrete lattices [26, 27, 28], leading to hallmark ground-state phenomena such as asymmetric momentum distributions [29, 30, 26, 31], Friedel oscillations [32], and statistics-induced phase transitions [27, 33, 34]. Experimentally, recent advances in ultracold atoms have brought the engineering of anyonic properties within reach [35, 36, 2, 38, 39, 40], making it both timely and essential to investigate the role of anyonic statistics in the nonequilibrium dynamics [41, 42, 43, 44, 45, 46] of quantum many-body systems.
@@
Far-from-equilibrium dynamics in bosonic and fermionic settings has attracted considerable attention in recent years [47, 48, 49, 50, 51, 52]. These quantum many-body systems exhibit rich emergent dynamical scaling behavior, providing valuable insights into collective quantum phenomena [53, 54, 55, 56, 57, 58]. Ballistic and diffusive transport have been observed experimentally in a broad range of quantum simulators [59, 60, 61, 62]. Beyond these conventional regimes, qualitatively new universality classes have also been identified, such as Kardar-Parisi-Zhang scaling in 1D quantum magnets [63, 64]. This behavior arises from the interplay of integrability and internal non-Abelian symmetries [65, 66, 67] and has been observed in quantum gas microscopy [68] and programmable quantum circuits [69]. These developments raise a fundamental question: can particle statistics itself give rise to universal scaling in far-from-equilibrium quantum dynamics beyond the Bose-Fermi paradigm?
@@
To address this question, we investigate the many-body relaxation dynamics of a 1D anyon-Hubbard model (AHM) [Fig. 1(a)], starting from far-from-equilibrium initial states. Upon varying the statistical angle $\theta$, we uncover emergent universal scaling behavior in the long-time dynamics that is distinct from conventional bosonic and fermionic limits. This feature manifests as a pronounced separation between the dynamics of density correlations and entanglement [Fig. 1(b)], reminiscent of particle-information decoupling discussed in many-body-localized systems [70, 71, 72, 73]. Specifically, the dynamical scaling of density correlations exhibits a crossover from ballistic ($z=1$) to superdiffusive ($z=3/2$) behavior as $\theta$ is varied. The anomalous scaling can be understood in terms of quantum interference between exchange paths of the holon-doublon pair, which suppresses its coherent propagation. In contrast, while density fluctuations spread superdiffusively, the von Neumann entanglement entropy grows ballistically. This growth is primarily carried by the configurational component, which propagates ballistically, leaving only a minor contribution from the number entropy. Finally, we show that this anomalous dynamical scaling is robust against onsite disorder and initially density configurations, offering a promising route for experimental verification in state-of-the-art quantum simulators.
@@
Figure 1: Anomalous far-from-equilibrium dynamical scaling in a 1D interacting anyonic system. (a) Schematic of a 1D anyonic lattice with hopping $J$, mapped onto a BHM with density-dependent tunneling $Je^{i\theta n}$ and interactions $U$, where $n$ and $\theta$ denote the local density and statistical angle, respectively. A quantum quench from an out-of-equilibrium state triggers the spreading of spatial correlations and entanglement throughout the system. (b) Distinct scaling of the half-chain particle-number fluctuations $\Delta N$ and the von Neumann entanglement entropy $S$. Both quantities are shown as functions of statistical angle $\theta$ and time $t$: for $U\leq J$, $\Delta N$ crosses over from ballistic to superdiffusive as $\theta$ increases from 0 to $\pi$, while $S$ remains ballistic.
@@
where $J$ and $U$ denote the nearest-neighbor hopping and onsite interaction strength, and $\hat{n}_{j}=\hat{a}_{j}^{\dagger}\hat{a}_{j}$ the anyon number operator on site $j$. The anyonic creation and annihilation operators $\hat{a}_{j}^{\dagger}$, $\hat{a}_{j}$ obey generalized commutation relations  [22, 74],
@@
$\hat{a}_{j}\hat{a}_{k}-e^{i\theta\text{sgn}(j-k)}\hat{a}_{k}\hat{a}_{j}=$ $0,$ (2b)
with statistical angle $\theta\in[0,\pi]$. The limits $\theta=0$ and $\theta=\pi$ recover bosonic and pseudofermionic statistics, respectively. An experimentally relevant route to simulate interacting anyons is to map the AHM onto a Bose-Hubbard model (BHM) with a density-dependent Peierls phase [75, 76, 77, 78], as illustrated in the top panel of Fig. 1(a). The mapped Hamiltonian reads [2]
@@
where $\hat{b}_{j}^{\dagger}$, $\hat{b}_{j}$ are bosonic operators related to the anyons via a generalized Jordan-Wigner transformation, $\hat{a}_{k}=\hat{b}_{k}e^{i\theta\sum_{j<k}\hat{n}_{j}}$. The statistical angle $\theta$ continuously interpolates between bosonic and pseudofermionic limits, reshaping the ground-state momentum distribution from a sharp condensate peak to a nearly flat fermionic profile [29, 30, 26, 31] and giving rise to rich anyonic dynamics—dynamical fermionization [41, 42], asymmetric transport [43, 44] and two-particle correlations [45]. Here we focus on the emergent universal scaling governing the far-from-equilibrium dynamics of interacting many-body anyonic systems.
@@
where $\hat{N}=\sum_{j=1}^{L/2}\hat{n}_{j}$ denotes the total particle number operator of the left half-chain. This quantity provides a probe of nonequilibrium particle transport and relaxation in lattice Hubbard models [62, 83, 84]. To quantify the spreading of quantum information, we further compute the von Neumann entanglement entropy,
@@
where $\rho_{A}(t)=\mathrm{Tr}_{B}[\lvert\psi(t)\rangle\langle\psi(t)\rvert$] is the reduced density matrix of the left half-chain $A$, obtained by tracing out its complement $B$. To extract the universal dynamical scaling of the observable dynamics, we rescale time according to $t\to t/L^{z}$ and define
@@
Anomalous scaling of density correlation— We first investigate the far-from-equilibrium scaling dynamics of density correlations in the anyonic chain for the noninteracting limit. We find that tuning the statistical angle $\theta$ drives a crossover from ballistic to superdiffusive spreading, in contrast to the ballistic dynamics of integrable bosonic and fermionic chains. This anomalous behavior is revealed by the CTD following a quench from the unit-filled product state $|\psi_{0}\rangle=|\cdots 111\cdots\rangle$. As shown in Fig. 2(a), following a short $\theta$-independent period, the characteristic length $l(t)$ exhibits robust power-law scaling, $l(t)\sim t^{\beta}$. The growth exponent $\beta$ decreases monotonically with increasing $\theta$, from $\beta=1$ in the bosonic limit ($\theta=0$) to $\beta=2/3$ at $\theta=\pi$ [inset of Fig. 2(a)]. In a 1D system, the corresponding dynamical exponent satisfies $z=1/\beta$ [90], establishing superdiffusive dynamics ($1<z<2$) induced by anyonic statistics, distinct from the bosonic and fermionic limits (z=1z=1[82, 90]. In finite systems, boundary effects can affect the scaling behavior. Here we find that a lattice size of $L=100$ already suffices to capture the anomalous scaling [90]. This scaling is independently confirmed by half-chain density fluctuations $\Delta N(t)$, which collapse onto a universal curve when time is rescaled as $t/L^{z}$, yielding $z=3/2$ at $\theta=\pi$ [Fig. 2(b)]. Although larger systems cannot be evolved long enough for $\Delta N(t)$ to reach the $L$-dependent saturation plateau, a small system of $L=14$ readily saturates, allowing accurate extraction of $z$ (inset of Fig. 2(b), [84]).
@@
We next examine the interplay between anyonic statistics and interactions by analyzing the CTD dynamics across a broad range of interaction strengths $U$. In the weakly interacting regime, the scaling exponent $\beta$ remains robust. As summarized in Fig. 2(d), the dynamics at $\theta=\pi$ is superdiffusive with $\beta\approx 2/3$ ($U\leq J$), while the bosonic limit ($\theta=0$) remains ballistic. By contrast, in the strong-interaction regime, the growth of $l(t)$ is nearly ballistic for all $\theta$, with $\beta\approx 1$, as shown in Fig. 2(c). This reflects the suppression of particle-exchange processes by strong interactions, rendering the effect of fractional statistics negligible and recovering conventional ballistic dynamics.
@@
The universal anomalous scaling is found to be robust against disorder and initial-state variations. We study $l(t)$ in the presence of finite onsite disorder and for different initial states, such as $|\cdots 110211\cdots\rangle$ and the density wave $|\cdots 0202\cdots\rangle$ [90]. In all cases, we find that the growth exponent $\beta$ remains largely unchanged, demonstrating that the anomalous scaling is intrinsically determined by anyonic statistics.
@@
Figure 3: Dynamical scaling of entanglement entropy in the relaxation dynamics of the anyonic chain described by $\hat{H}$. (a) Rescaled von Neumann entropy $\widetilde{S}(t)$ as a function of time $t$ for $\theta=\pi/2$, with the inset showing the unrescaled data. The nearly ballistic scaling with $z\simeq 1$ demonstrates a clear separation between entanglement propagation and particle transport [Fig. 2(d)]. (b) Schematic illustration of the reduced density matrix $\rho_{A}$, where the total entanglement entropy $S$ is decomposed into the number entropy $S_{N}$, associated with particle-number fluctuations between sectors, and the configuration entropy $S_{C}$, arising from superpositions within each particle-number sector. (c) Number entropy $\widetilde{S}_{N}(t)$ and (d) configuration entropy $\widetilde{S}_{C}(t)$ for $\theta=\pi/2$. Throughout the evolution, the entanglement growth is dominated by $S_{C}$ (insets). Results are obtained for system sizes L=10L=101414 at interaction strength $U/J=0.2$, using bond dimensions up to 4000.
@@
Ballistic dynamical scaling in entropy— We now turn to the dynamics of the entanglement entropy S(t){S}(t). While information propagation is generally expected to follow the identical temporal scaling behavior as particle transport [92, 93, 94], we surprisingly find that the entropy grows ballistically and is essentially independent of the statistical angle $\theta$. Focusing on the anyonic case $\theta=\pi/2$ in the weakly interacting regime ($U/J=0.2$) and starting from the product state $|\psi_{0}\rangle=|\cdots 111\cdots\rangle$, $S(t)$ as expected grows from zero to an $L$-dependent saturation value, as shown in the inset of Fig. 3(a). Remarkably, data for different system sizes collapse when time is rescaled as $t\rightarrow t/L^{z}$, yielding $z\approx 1.05$ (see Fig. 3(a) and Ref. [90]). This nearly ballistic growth contrasts with the anomalous scaling of density fluctuations $\Delta N(t)$ [Fig. 2(d), $z\approx 1.62$], revealing a clear separation between quantum-information propagation and particle transport. Unlike many-body localized systems, where such separation is induced by disorder [70, 71, 72, 73], here it emerges intrinsically from anyonic statistics.
@@
Figure 4: Diagram of interference picture of holon-doublon pairs under anyonic statistics and experimental realization. (a) Schematic illustration of holon-doublon propagation starting from an initially adjacent holon-doublon pair ($d=1$) to separation $d=3$ via two distinct hopping paths. Paths along these trajectories acquire $\theta$-dependent phases, leading to destructive interference that suppresses the propagation of the pairs. (b) Density deviation dynamics $\delta n_{j}(t)$ governed by Eq. (3), after a central-site excitation for the bosonic (left, $\theta=0$), anyonic (middle, $\theta=\pi/2$), and pseudofermionic (right, $\theta=\pi$) cases. (c) Floquet protocol realized in Ref. [2]: three-tone modulation $V=\sum V_{i}$ (solid) resonantly restores tunneling in a tilted lattice; $V_{1,2,3}$ (dashed, dotted, dot-dashed) selectively drive singly-to-empty, singly-to-singly, and doubly-to-empty transitions. The phase offset $\theta$ of $V_{3}$ implements the density-dependent Peierls phase. (d) Density deviation dynamics $\delta n_{j}(t)$ for the setup in (c). The agreement demonstrates the robust realization of anyonic dynamics in the many-body regime. Time $t$ is in units of $1/J$ in (b) and $1/J^{\prime}$ in (d), with $J^{\prime}$ the effective Floquet hopping rate [90]. Results are obtained for a system of size $L=100$ in the weakly interacting regime ($U/J=0.2$).
@@
Experimental accessibility— The identified many-body features are within reach of current ultracold atom experiments, leveraging the Floquet protocol recently demonstrated in Ref. [2, 40]. While this protocol was originally demonstrated in a two-particle setting, the underlying interference mechanism remains robust in the many-body regime. As detailed in Fig. 4(c), a three-tone modulation of the lattice depth resonantly restores tunneling otherwise suppressed by the static tilt. Crucially, the phase offset $\theta$ of the $V_{3}$ component (dot-dashed line) implements the density-dependent Peierls phase. The numerical results in Fig. 4(d) exhibit agreement between the full Floquet-driven dynamics and the direct evolution of Eq. (3). Note that the difference at $\theta=\pi/2$ originates from the opposite statistical angle acquired during the Floquet realization in a many-body setting [90]. Despite such a discrepancy, the overall dynamics exhibits qualitative agreement. This indicates that the additional complexities inherent in the Floquet scheme—such as higher-order tunneling—would not alter the essential physics, thereby establishing the predicted phenomena as an experimentally accessible feature. Furthermore, single-site-resolved quantum-gas microscopy provides direct access to both correlation functions [99, 93, 100, 84] and entanglement entropy [101, 102, 95], offering a powerful experimental probe into nonequilibrium dynamics governed by fractional statistics.
@@
Conclusion—We have studied the far-from-equilibrium quench dynamics of interacting anyons in a 1D lattice, revealing unique dynamical scaling governed by anyonic statistics. In the weakly interacting regime, particle-number fluctuations cross over from ballistic to superdiffusive, a behavior beyond both bosonic and fermionic limits. Remarkably, while number fluctuations spread superdiffusively, the half-chain entanglement entropy remains ballistic, demonstrating a clear separation between particle and information propagation. This anomalous scaling is connected to quantum interference effects induced by the anyonic statistical phase, which suppress the propagation of coherent holon-doublon pairs. Our results provide new insights into nonequilibrium physics governed by anyonic statistics and advance the control of transport and entanglement in synthetic quantum matter, realizable with current experimental techniques [2]. Future work may clarify the universality classes underlying the predicted scaling using high-order cumulants in the long-time dynamics of larger quantum systems [68, 69]. It is an interesting question to explore richer nonequilibrium dynamics and entanglement growth beyond one dimension [5, 15].
@@
We acknowledge useful discussions with Yinghai Wu and Felix Helluin. This work is supported by the National Natural Science Foundation of China (Grants No. 12374252, No. 12074431, No. 12174130, No. 12304076, and No. 12404332), and the Excellent Youth Foundation of Hunan Scientific Committee under Grant No. 2021JJ10044. Numerical calculations were performed on the TianHe-1A cluster of the National Supercomputer Center at Tianjin, and HPC resources on the ChinaHPC.
@@
“Anomalous dynamical scaling in interacting anyonic chains”
@@
$i\frac{d}{dt}\hat{b}_{i}(t)=[\hat{b}_{i}(t),\hat{H}_{0}]=-J\left[\hat{b}_{i-1}(t)+\hat{b}_{i+1}(t)\right].$ (S2)
@@
As an additional check, we simulate dynamics from a half-filling initial state $|\cdots 0101\cdots\rangle$, shown in Fig. S3(g)(h). Due to the limited accessible evolution time, the dynamical scaling of $l(t)$ appears nearly ballistic in both the weakly and strongly interacting regimes. These results provide an important complement to the dynamical scaling analysis of the anyonic Hubbard chain. How the anomalous scaling depends on the filling factor remains to be explored in future work.
Figure S3: Robustness of the anomalous dynamical scaling. (a–d) Anomalous dynamical scaling starting from different initial states. (e,f) Robustness to the onsite disorder term $H_{\mathrm{dis}}$. For finite disorder strength $W=0.5$, $l(t)$ remains superdiffusive. (g,h) Dynamical scaling of $l(t)$ starting from half-filling initial state $|\cdots 0101\cdots\rangle$.

### 22 Mar 2026 v2

# Statistics-governed dynamical scaling in interacting anyonic chains

@@ Abstract
Particle statistics impose fundamental constraints on nonequilibrium quantum dynamics, yet it remains an open question whether anyonic statistics can lead to emergent dynamical scaling beyond the conventional Bose-Fermi paradigm. Here we investigate the far-from-equilibrium many-body relaxation of anyons in a one-dimensional lattice, uncovering a statistics-governed, robust scaling behavior that deviates from standard Bose–Fermi limits. Based on large-scale numerical simulations and scaling analysis, we find that in the weakly interacting regime, anyonic statistics leads to emergent superdiffusive scaling in particle transport, while the entanglement entropy remains ballistic and is essentially insensitive to exchange statistics. The anomalous dynamics can be interpreted intuitively from the statistical-phase-induced quantum interference that suppresses coherent holon-doublon propagation; in contrast, the entanglement growth is dominated by its configurational component, which maintains ballistic spreading regardless of the statistical phase. Our results establish anyonic statistics as a distinct source of universal nonequilibrium dynamics beyond bosons and fermions, with direct relevance to current quantum simulation experiments.
@@
Introduction— Anyonic statistics provides a fundamental extension of quantum many-body physics beyond the conventional Bose-Fermi framework [1, 2, 3, 4, 5]. Its physical relevance has been established through the observation of exotic quantum phenomena, including the fractional quantum Hall effect [6, 7, 8, 9] and topological spin liquids [10, 11, 12], and has stimulated extensive interest in topological quantum computation [13, 14, 15, 16]. While anyonic statistics are intrinsically associated with two-dimensional systems, anyonic behavior can also emerge in one dimension (1D). For instance, fractional exclusion statistics were first revealed through the generalized Pauli principle in the Haldane-Shastry model [17, 18, 19]. This has triggered extensive studies of 1D anyonic models both in the continuum [20, 21, 22] and on discrete lattices [23, 24, 25, 26], leading to hallmark ground-state phenomena such as asymmetric momentum distributions [27, 28, 23, 29], Friedel oscillations [30], and statistics-induced phase transitions [24, 31, 32]. Experimentally, recent advances in ultracold atoms have brought the engineering of anyonic properties within reach [33, 34, 35, 36, 37], making it both timely and essential to investigate the role of anyonic statistics in the nonequilibrium dynamics [38, 39, 40, 41, 42, 43] of quantum many-body systems.
@@
Far-from-equilibrium dynamics in bosonic and fermionic settings has attracted considerable attention in recent years [44, 45, 46, 47, 48, 49]. These quantum many-body systems exhibit rich emergent dynamical scaling behavior, providing valuable insights into collective quantum phenomena [50, 51, 52, 53, 54, 55]. Ballistic and diffusive transport have been observed experimentally in a broad range of quantum simulators [56, 57, 58, 59]. Beyond these conventional regimes, qualitatively new universality classes have also been identified, such as Kardar-Parisi-Zhang scaling in 1D quantum magnets [60, 61]. This behavior arises from the interplay of integrability and internal non-Abelian symmetries [62, 63, 64] and has been observed in quantum gas microscopy [65] and programmable quantum circuits [66]. These developments raise a fundamental question: can particle statistics itself give rise to emergent scaling in far-from-equilibrium quantum dynamics beyond the Bose-Fermi paradigm?
@@
To address this question, we investigate the many-body relaxation dynamics of a 1D anyon-Hubbard model (AHM) [Fig. 1(a)], starting from far-from-equilibrium initial states. We demonstrate that anyonic statistics can induce emergent dynamical scaling that goes beyond conventional bosonic and fermionic limits. This effect is particularly prominent in the weak-interaction regime: while the density correlations in these conventional limits exhibit ballistic transport, those in anyonic systems undergo a distinct crossover to superdiffusive behavior as the statistical angle $\theta$ is varied [Fig. 1(b)]. The anomalous scaling can be understood in terms of quantum interference between exchange paths of the holon-doublon pair, which suppresses its coherent propagation. While anyonic interference suppresses particle transport, the entanglement entropy remains robustly ballistic. This contrast arises because the configuration entropy dominates the growth and propagates ballistically, whereas the number entropy is primarily constrained by the statistical phase. Finally, we show that this anomalous dynamical scaling is robust against onsite disorder and initially density configurations, offering a promising route for experimental verification in state-of-the-art quantum simulators.
@@
Figure 1: Emergent far-from-equilibrium dynamical scaling in a 1D interacting anyonic system. (a) Schematic of a 1D anyonic lattice with hopping $J$, mapped onto a BHM with density-dependent tunneling $Je^{i\theta n}$ and interactions $U$, where $n$ and $\theta$ denote the local density and statistical angle, respectively. A quantum quench from an out-of-equilibrium state triggers the spreading of spatial correlations and entanglement throughout the system. (b) Distinct scaling of the half-chain particle-number fluctuations $\Delta N$ and the von Neumann entanglement entropy $S$. Both quantities are shown as functions of statistical angle $\theta$ and time $t$: for $U\leq J$, $\Delta N$ crosses over from ballistic to superdiffusive as $\theta$ increases from 0 to $\pi$, while $S$ remains ballistic.
@@
where $J$ denotes the nearest-neighbor hopping amplitude, $U$ onsite interaction strength, and $\hat{n}_{j}=\hat{a}_{j}^{\dagger}\hat{a}_{j}$ the anyon number operator on site $j$. The anyonic creation and annihilation operators $\hat{a}_{j}^{\dagger}$, $\hat{a}_{j}$ obey generalized commutation relations  [20, 67],
@@
$\hat{a}_{j}\hat{a}_{k}-e^{i\theta\text{sgn}(j-k)}\hat{a}_{k}\hat{a}_{j}=$ $0.$ (2b)
An experimentally relevant route to simulate interacting anyons is to map the AHM onto a Bose-Hubbard model (BHM) with a density-dependent Peierls phase [68, 69, 70, 71], as illustrated in the top panel of Fig. 1(a). The mapped Hamiltonian reads [35, 37]
@@
where $\hat{b}_{j}^{\dagger}$, $\hat{b}_{j}$ are bosonic operators related to the anyons via a generalized Jordan-Wigner transformation, $\hat{a}_{k}=\hat{b}_{k}e^{i\theta\sum_{j<k}\hat{n}_{j}}$. The statistical angle $\theta$ continuously interpolates between bosonic and pseudofermionic limits, reshaping the ground-state momentum distribution from a sharp condensate peak to a nearly flat fermionic profile [27, 28, 23, 29] and giving rise to rich anyonic dynamics, including dynamical fermionization [38, 41], asymmetric transport [42, 43], and two-particle correlations [39]. Here we focus on the emergent scaling governing the far-from-equilibrium dynamics of interacting many-body anyonic systems.
@@
where $\hat{N}=\sum_{j=1}^{L/2}\hat{n}_{j}$ denotes the total particle number operator of the left half-chain. This quantity provides a probe of nonequilibrium particle transport and relaxation in lattice Hubbard models [59, 76, 77]. To quantify the spreading of quantum information, we compute the von Neumann entanglement entropy,
@@
where $\rho_{A}(t)=\mathrm{Tr}_{B}[\lvert\psi(t)\rangle\langle\psi(t)\rvert$] is the reduced density matrix of the left half-chain $A$, obtained by tracing out its complement $B$. To extract the dynamical scaling of the observable dynamics, we rescale time according to $t\to t/L^{z}$ and define
@@
Anomalous scaling of density correlation— We first investigate the far-from-equilibrium dynamics of density correlations in the anyonic chain in the noninteracting limit. We find that tuning the statistical angle $\theta$ drives a crossover from ballistic to superdiffusive spreading, in contrast to the ballistic dynamics of integrable bosonic and fermionic chains. This anomalous behavior is revealed by the CTD following a quench from the unit-filled product state $|\psi_{0}\rangle=|\cdots 111\cdots\rangle$. As shown in Fig. 2(a), following a short $\theta$-independent period, the characteristic length $l(t)$ exhibits robust power-law scaling, $l(t)\sim t^{\beta}$. The growth exponent $\beta$ decreases monotonically with increasing $\theta$, from $\beta=1$ in the bosonic limit ($\theta=0$) to $\beta=2/3$ at $\theta=\pi$ [inset of Fig. 2(a)]. In a 1D system, the corresponding dynamical exponent satisfies $z=1/\beta$ [83], establishing superdiffusive dynamics ($1<z<2$) induced by anyonic statistics, distinct from the bosonic and fermionic limits (z=1z=1[84, 75, 83]. In finite systems, boundary effects can affect the scaling behavior. Here we find that a lattice size of $L=100$ already suffices to capture the anomalous scaling [83]. This scaling is independently confirmed by half-chain density fluctuations $\Delta N(t)$, which collapse onto a universal curve when time is rescaled as $t/L^{z}$, yielding $z=3/2$ at $\theta=\pi$ [Fig. 2(b)]. Although larger systems cannot be evolved long enough for $\Delta N(t)$ to reach the $L$-dependent saturation plateau, a small system of $L=14$ readily saturates, allowing accurate extraction of $z$ (inset of Fig. 2(b), [77]).
@@
We next examine the interplay between anyonic statistics and interactions by analyzing the CTD dynamics across a broad range of interaction strengths $U$. In the weakly interacting regime, the scaling exponent $\beta$ remains robust. As summarized in Fig. 2(d), the dynamics at $\theta=\pi$ is superdiffusive with $\beta\approx 2/3$ ($U\leq J$), while the bosonic limit ($\theta=0$) remains ballistic. By contrast, in the strong-interaction regime, the growth of $l(t)$ is nearly ballistic for all $\theta$, with $\beta\approx 1$, as shown in Fig. 2(c). This reflects the suppression of particle-exchange processes by strong interactions, rendering the effect of anyonic statistics negligible and recovering conventional ballistic dynamics.
@@
The anomalous scaling is found to be robust against disorder and initial-state variations. We study $l(t)$ in the presence of finite onsite disorder and for different initial states, such as $|\cdots 110211\cdots\rangle$ and the density wave $|\cdots 0202\cdots\rangle$ [83]. In all cases, we find that the growth exponent $\beta$ remains largely unchanged, demonstrating that the anomalous scaling is intrinsically determined by anyonic statistics.
@@
Figure 3: Dynamical scaling of entanglement entropy in the relaxation dynamics of the anyonic chain. (a) Rescaled von Neumann entropy $\widetilde{S}(t)$ as a function of time $t$ for $\theta=\pi/2$, with the inset showing the unrescaled data. The nearly ballistic scaling with $z\simeq 1$ demonstrates a clear separation between entanglement propagation and particle transport [Fig. 2(d)]. (b) Schematic illustration of the reduced density matrix $\rho_{A}$, where the total entanglement entropy $S$ is decomposed into the number entropy $S_{N}$, associated with particle-number fluctuations between sectors, and the configuration entropy $S_{C}$, arising from superpositions within each particle-number sector. (c) Number entropy $\widetilde{S}_{N}(t)$ and (d) configuration entropy $\widetilde{S}_{C}(t)$ for $\theta=\pi/2$. Throughout the evolution, the entanglement growth is dominated by $S_{C}$ (insets). Results are obtained for system sizes L=10L=101414 at interaction strength $U/J=0.2$, using bond dimensions up to 4000.
@@
Ballistic dynamical scaling in entropy— We now turn to the dynamics of the entanglement entropy S(t){S}(t). While information propagation typically follows the same ballistic scaling as particle transport in the noninteracting bosonic [76] and fermionic limits [85], we find that in anyonic chains, the entropy remains ballistic and essentially independent of the statistical angle $\theta$, despite the significant modification of the particle dynamics. Focusing on an anyonic case with $\theta=\pi/2$ in the weakly interacting regime ($U/J=0.2$) and starting from the product state $|\psi_{0}\rangle$, $S(t)$ as expected grows from zero to an $L$-dependent saturation value, as shown in the inset of Fig. 3(a). Remarkably, data for different system sizes collapse when time is rescaled as $t\rightarrow t/L^{z}$, yielding $z\approx 1.05$ (see Fig. 3(a) and Ref. [83]). This ballistic growth contrasts with the anomalous scaling of density fluctuations $\Delta N(t)$ [Fig. 2(d), $z\approx 1.62$], highlighting the different dynamical scales associated with information and particle spreading. Unlike localized systems where both particle and information spreading are simultaneously suppressed by disorder [86, 87, 88], the phenomenon predicted here is driven by anyonic statistics.
@@
Figure 4: Diagram of interference picture of holon-doublon pairs under anyonic statistics and experimental realization. (a) Schematic illustration of holon-doublon propagation starting from an initially adjacent holon-doublon pair ($d=1$) to separation $d=3$ via two distinct hopping paths. Paths along these trajectories acquire $\theta$-dependent phases, leading to destructive interference that suppresses the propagation of the pairs. (b) Density deviation dynamics $\delta n_{j}(t)$ governed by Eq. (3), after a central-site excitation for the bosonic (left, $\theta=0$), intermediate (middle, $\theta=\pi/2$), and pseudofermionic (right, $\theta=\pi$) cases. Dashed lines indicate the propagation front of the pair. (c) Floquet protocol realized in Ref. [35]: three-tone modulation $V=\sum V_{i}$ (solid) resonantly restores tunneling in a tilted lattice; $V_{1,2,3}$ (dashed, dotted, dot-dashed) selectively drive singly-to-empty, singly-to-singly, and doubly-to-empty transitions. The phase offset $\theta$ of $V_{3}$ implements the density-dependent Peierls phase. (d) Density deviation dynamics $\delta n_{j}(t)$ for the setup in (c). The agreement demonstrates the robust realization of anyonic dynamics in the many-body regime. Time $t$ is in units of $1/J$ in (b) and $1/J^{\prime}$ in (d), with $J^{\prime}$ the effective hopping strength [83]. Results are obtained for a system of size $L=100$ in the weakly interacting regime ($U/J=0.2$).
@@
Experimental accessibility— The identified many-body features are within reach of current ultracold atom experiments, leveraging the Floquet protocol recently demonstrated in Ref. [35, 37]. While this protocol was originally demonstrated in a two-particle setting, the underlying interference mechanism remains robust in the many-body regime. As detailed in Fig. 4(c), a three-tone modulation of the lattice depth resonantly restores tunneling otherwise suppressed by the static tilt. Crucially, the phase offset $\theta$ of the $V_{3}$ component (dot-dashed line) implements the density-dependent Peierls phase. The numerical results in Fig. 4(d) exhibit agreement between the full Floquet-driven dynamics and the direct evolution of Eq. (3). Note that the difference at $\theta=\pi/2$ originates from the opposite statistical angle acquired during the Floquet realization in a many-body setting [83]. Despite such a discrepancy, the overall dynamics exhibits qualitative agreement. This indicates that the additional complexities inherent in the Floquet scheme, such as higher-order tunneling, would not alter the essential physics, thereby establishing the predicted phenomena as an experimentally accessible feature. Furthermore, single-site-resolved quantum-gas microscopy provides direct access to both correlation functions [94, 93, 95, 77] and entanglement entropy [96, 97, 89], offering a powerful experimental probe into nonequilibrium dynamics governed by anyonic statistics.
@@
Conclusion— In summary, we have demonstrated that anyonic statistics fundamentally modifies far-from-equilibrium quantum dynamics, driving a robust dynamical scaling that extends beyond the conventional bosonic and fermionic limits. In contrast to previous studies focusing on asymmetric transport [42, 43, 35, 37] or dynamical fermionization [38, 41], our work identifies a distinct regime of scaling dynamics: while particle transport enters a superdiffusive regime, the entanglement entropy growth remains ballistic. This anomalous scaling is connected to quantum interference effects induced by the anyonic statistical phase, which suppress the propagation of coherent holon-doublon pairs. The scaling exponents exhibit a weak dependence on initial states within the explored regime, indicating that this statistics-induced dynamical scaling is a characteristic feature of anyonic systems. Our results provide new insights into nonequilibrium physics governed by anyonic statistics and advance the control of transport and entanglement in synthetic quantum matter, realizable with current experimental techniques [35, 37]. Future work may clarify the universality classes underlying the predicted scaling using high-order cumulants in the long-time dynamics of larger quantum systems [65, 66]. It is an interesting question to explore richer nonequilibrium dynamics and entanglement growth beyond one dimension [15, 98, 99, 5, 100].
@@
We acknowledge useful discussions with Yinghai Wu, and Felix Helluin. This work is supported by the National Natural Science Foundation of China (Grants No. 12374252, No. 12074431, No. 12174130, No. 12304076, and No. 12404332), and the Excellent Youth Foundation of Hunan Scientific Committee under Grant No. 2021JJ10044. Numerical calculations were performed on the TianHe-1A cluster of the National Supercomputer Center at Tianjin, and HPC resources on the ChinaHPC.
@@
“Statistics-governed dynamical scaling in interacting anyonic chains”
@@
$i\frac{d}{dt}\hat{b}_{j}(t)=[\hat{b}_{j}(t),\hat{H}_{0}]=-J\left[\hat{b}_{j-1}(t)+\hat{b}_{j+1}(t)\right].$ (S2)
@@
Figure S3: Robustness of the anomalous dynamical scaling. (a–d) Anomalous dynamical scaling starting from different initial states. (e,f) Robustness to the onsite disorder term $H_{\mathrm{dis}}$. For finite disorder strength $W=0.5$, $l(t)$ remains superdiffusive.